# ThinkingViT: Matryoshka Thinking Vision Transformer for Elastic Inference

### v2

@@ 4.3 Comparing ThinkingViT and baselines
Figure 1 compares ThinkingViT with SOTA baselines in terms of GMACs and inference throughput, using the 3H $\rightarrow$ 6H configuration. ThinkingViT achieves up to 2.0 p.p. higher accuracy at equal throughput, and up to 2.9 p.p. higher accuracy at the same GMACs compared with its baselines. This improvement stems from three key factors. First, ThinkingViT makes early predictions using only 3 heads for easy inputs and expands to 6 heads only when needed. Second, ThinkingViT fuses embeddings from the first stage into the second, leading to better performance than flat 6-head models like HydraViT or DeiT-S. Moreover, ThinkingViT achieves 81.44% with only 22.01 M params, which is just 0.36 p.p. lower than DeiT-Base with 86.6 M params. This result further validates the effectiveness of the Token Recycling strategy. Importantly, it also demonstrates that achieving higher accuracy does not necessarily require a larger model; rather, it depends on applying sufficient computation. Third, baselines typically employ multiple slicing points throughout the architecture to meet diverse GMACs. However, optimizing for all such configurations can lead to accuracy reduction [haberer2024hydravit]. ThinkingViT sidesteps this issue by supporting elastic inference using just a few nested subnetworks, with compute budget adaptability controlled via the entropy threshold. We report detailed performance results in the Appendix A.
@@ 4.5 Semantic Segmentation
ThinkingViT builds on the vanilla ViT architecture and is therefore compatible with models designed for ViT backbones. To verify this, we replace the ViT backbone in the Segmenter [strudel2021segmenter] model with ThinkingViT and employ a linear decoder for semantic segmentation. We evaluate on the ADE20K [zhou2017scene] and Cityscapes [cordts2016cityscapes] datasets using both single-scale and multi-scale inference. As shown in Fig. 9, ThinkingViT consistently outperforms DeiT-Small and DeiT-Tiny backbones, demonstrating its effectiveness as a drop-in replacement for vision transformer architectures in downstream tasks. Furthermore, Fig. 10 visualizes how ThinkingViT refines object boundaries and improves segmentation accuracy with additional rounds of "thinking". ThinkingViT also extends to other ViT based tasks such as DETR [carion2020end] style object detection, which we leave for future work.
@@ 4.6 Extension to Hierarchical Architectures
While all previous experiments were based on a vanilla ViT architecture, we also demonstrate that ThinkingViT can be applied to more modern variants of Vision Transformers, such as the Swin Transformer [liu2021swin]. The key difference in Swin Transformer is that it consists of four stages, each with an increasing number of attention heads and a decreasing number of tokens, which grow progressively larger. We integrate ThinkingViT on top of Swin-S, by allocating the entire first stage and half of the capacity of the last three stages for the first round, as these contain most of the layers and attention heads. The second round is then executed using the full model. As illustrated in Figure 8(a), this design allows ThinkingViT with a Swin backbone to achieve the same accuracy as Swin-B (83.5%) while requiring only 9.5 GMACs and 50M parameters, compared to 15.4 GMACs and 88M parameters for Swin-B. Moreover, our model achieves a higher accuracy per GMAC compared to both Swin-S and Swin-T, demonstrating that our approach can improve performance even for hierarchical architectures.
@@ 5 Conclusion
We introduce ThinkingViT, a nested ViT that uses progressive thinking stages to adapt inference computation to input difficulty. The model begins with a small subset of important attention heads and halts early when confidence is high. Otherwise, within the same backbone, it activates a larger subset of attention heads and conducts a new forward pass. This iterative expansion continues until the model reaches the desired confidence level or exhausts its maximum capacity. Token Recycling further improves later stages by fusing input embeddings with features from previous stages. ThinkingViT outperforms both width-based and depth-based expansion strategies, and its backbone-preserving design enables seamless transfer to downstream tasks such as semantic segmentation. We also demonstrate that it extends effectively to architectures such as Swin.
@@ Appendix B Results on ImageNet variants
Figure 12 shows the performance of ThinkingViT on ImageNet variants. Note that the GMACs for ImageNet-R and ImageNet-V2 are reported in Figure 4(c) and Figure 4(b), respectively.
@@ Appendix D Ablation of different token recycling strategies
In Table 4, we compare several design variants for recycling information from the first round of inference to get better performance in the second round. We conduct these experiments on ThinkingViT 3H $\rightarrow$ 6H. In the Layerwise Activation Snapshots variant, we cache the hidden states at each of the 12 layers from the first round and feed them into the corresponding layers in the second round. However, this approach performs suboptimally, likely because it forces all intermediate representations from the first round to be reused in the second round, which reduces the performance. In the Memory Tokens design, we introduce a set of learnable memory tokens [darcet2024vision] that, in the first round, store helpful information for the second round. We also experiment with hybrid strategies that combine activation snapshots with Memory Tokens or introduce a fresh [CLS] token in the second round. We additionally evaluated the use of KV-cache reuse [vaswani2017attention], originally introduced for generative models, and include the results in Table 4. Ultimately, we find that the simplest strategy, Final-Layer Token Recycling, which reuses the features from the last layer of the first round, achieves the best performance. This indicates that the output of the first round already captures sufficient high-level information to guide the second round effectively, while being comparatively simple to implement. It is important to note that except the KV-cache experiment, all results were conducted during the early experimental phase of ThinkingViT, and due to training resource constraints, they were not trained with the full joint training strategy described in Section 3.3, but rather with the stochastic training method introduced in [haberer2024hydravit]. As a result, their accuracies are slightly lower than those of the final model reported in Section 4.
@@ Appendix E Ablation of different fusing strategies
Table 5 compares six fusion strategies that recycle the tokens produced in the first round, denoted as $z_{1}$, by incorporating them into the initial patch embeddings of the second round, $\mathcal{E}_{2}$, on ThinkingViT with a configuration of $3H\rightarrow 6H$. Each variant forms the second-round input as a linear blend, where $\alpha$ is a learnable scalar. The strategies differ in how the two embeddings are projected to the same dimensional space. We consider two projection strategies: (I) a parameter-free approach that either repeats the embedding or pads the lower-dimensional embeddings with zeros, and (II) a learnable linear projection. Empirically, the learnable projection achieves the best overall accuracy by having the highest accuracy in the second round while maintaining performance comparable to the strongest model configuration in the first round. In addition, we evaluate different initializations of $\alpha$ and observe that initializing with $\alpha=0$ consistently leads to the best performance. This initialization allows the model to begin training without relying on information from the first round, thereby enabling it to learn the optimal degree of reuse. For example, in ThinkingViT configured with 3H $\rightarrow$ 6H heads, the learned value of $\alpha$ converges to 0.19-0.19.
@@ Appendix G Comparing and combining ThinkingViT with other adaptive baselines
We note that ThinkingViT performs routing at the image level, which is orthogonal to token-level pruning, and thus can be combined with such approaches to provide additional flexibility. To demonstrate this, we incorporate the token pruning mechanism of DynamicViT [rao2021dynamicvit] into the second round of ThinkingViT 3H $\rightarrow$ 6H, applying a pruning ratio of 0.8. As shown in Fig. 14, integrating DynamicViT further improves both efficiency and accuracy, indicating that ThinkingViT 3H $\rightarrow$ 6H is compatible with and complementary to existing token pruning approaches.
@@
Appendix I Limitations
Training Overhead: Compared to training multiple standalone models, nested models like ThinkingViT incur higher training costs to reach similar performance levels across all stages. This is a known limitation of nested architectures, which share parameters and require joint optimization to maintain accuracy at varying compute levels.
Jumping Too Far Reduces Effectiveness: When the model transitions from a very small subset of attention heads (e.g., 3H) directly to a full configuration (e.g., 12H) in the second stage, performance suffers compared to using an intermediate stage (e.g., 9H) instead. This suggests that overly aggressive compute expansion can disrupt representational continuity, emphasizing the importance of smooth progression in staged inference.
Appendix J Batch inference
@@ Appendix I Limitations
Appendix K Effect of attention-head expansion and loss weighting on multi-round thinking

### v3

@@ 4.3 Comparing ThinkingViT and baselines
Figure 1 compares ThinkingViT with SOTA baselines in terms of GMACs and inference throughput, using the 3H $\rightarrow$ 6H configuration. ThinkingViT achieves up to 2.0 p.p. higher accuracy at equal throughput, and up to 2.9 p.p. higher accuracy at the same GMACs compared with its baselines. This improvement stems from three key factors. First, ThinkingViT makes early predictions using only 3 heads for easy inputs and expands to 6 heads only when needed. Second, ThinkingViT fuses embeddings from the first stage into the second, leading to better performance than flat 6-head models like HydraViT or DeiT-S. Moreover, ThinkingViT achieves 81.44% with only 22.01 M parameters, which is just 0.36 p.p. lower than DeiT-Base with 86.6 M parameters. This result further validates the effectiveness of the Token Recycling strategy. Importantly, it also demonstrates that achieving higher accuracy does not necessarily require a larger model; rather, it depends on applying sufficient computation. Third, baselines typically employ multiple slicing points throughout the architecture to meet diverse GMACs. However, optimizing for all such configurations can lead to accuracy reduction [haberer2024hydravit]. ThinkingViT sidesteps this issue by supporting elastic inference using just a few nested subnetworks, with compute budget adaptability controlled via the entropy threshold. We report detailed performance results in the Appendix A.
@@ 4.5 Semantic Segmentation
ThinkingViT builds on the vanilla ViT architecture and is therefore compatible with models designed for ViT backbones. To verify this, we replace the ViT backbone in the Segmenter [strudel2021segmenter] model with ThinkingViT and employ a linear decoder for semantic segmentation. We evaluate on the ADE20K [zhou2017scene] and Cityscapes [cordts2016cityscapes] datasets using both single-scale and multi-scale inference. As shown in Fig. 9, ThinkingViT consistently outperforms DeiT-Small and DeiT-Tiny backbones, demonstrating its effectiveness as a drop-in replacement for vision transformer architectures in downstream tasks. Furthermore, Fig. 10 visualizes how ThinkingViT refines object boundaries and improves segmentation accuracy with additional rounds of "thinking". ThinkingViT is also compatible with ViT based tasks such as DETR [carion2020end] style object detection, which we leave for future work.
@@ 4.6 Extension to Hierarchical Architectures
While all previous experiments were based on a vanilla ViT architecture, we also demonstrate that ThinkingViT can be applied to more modern variants of Vision Transformers, such as the Swin Transformer [liu2021swin]. The key difference in Swin Transformer is that it consists of four stages, each with an increasing number of attention heads and a decreasing number of tokens, which grow progressively larger. We integrate ThinkingViT on top of Swin-S, by allocating the entire first stage and half of the capacity of the last three stages for the first round, as the latter contains most of the layers and attention heads. The second round is then executed using the full model. As illustrated in Figure 8(a), this design allows ThinkingViT with a Swin backbone to achieve the same accuracy as Swin-B (83.5%) while requiring only 9.5 GMACs and 50M parameters, compared to 15.4 GMACs and 88M parameters for Swin-B. Moreover, our model achieves a higher accuracy per GMAC compared to both Swin-S and Swin-T, demonstrating that our approach can improve performance even for hierarchical architectures.
@@ 5 Conclusion
We introduce ThinkingViT, a nested ViT that uses progressive thinking stages to adapt inference computation to input difficulty. The model begins with a small subset of important attention heads and halts early when confidence is high. Otherwise, within the same backbone, it activates a larger subset of attention heads and conducts a new forward pass. This iterative expansion continues until the model reaches the desired confidence level or exhausts its maximum capacity. Token Recycling further improves later stages by fusing input embeddings with features from previous stages. ThinkingViT outperforms both width-based and depth-based expansion strategies, and its backbone-preserving design enables seamless transfer to downstream tasks such as semantic segmentation. Experiments show that ThinkingViT surpasses nested baselines by up to 2.0 percentage points (p.p.) in accuracy at the same throughput and by up to 2.9 p.p. at equal GMACs on ImageNet-1K. We also demonstrate that it extends effectively to architectures such as Swin Transformer.
@@ Appendix B Results on ImageNet variants
Figure 11 and  12 show the performance of ThinkingViT on ImageNet variants. Note that the GMACs for ImageNet-V2 and ImageNet-R are reported in Figure 4(b) and Figure 4(c), respectively.
@@ Appendix D Ablation of different token recycling strategies
In Table 4, we compare several design variants for recycling information from the first round of inference to get better performance in the second round. We conduct these experiments on ThinkingViT 3H $\rightarrow$ 6H. In the Layerwise Activation Snapshots variant, we cache the hidden states at each of the 12 layers from the first round and feed them into the corresponding layers in the second round. However, this approach performs suboptimally, likely because it forces all intermediate representations from the first round to be reused in the second round, which makes it harder for the tokens as they need to work for both rounds. In the Memory Tokens design, we introduce a set of learnable memory tokens [darcet2024vision] that, in the first round, store helpful information for the second round. We also experiment with hybrid strategies that combine activation snapshots with Memory Tokens or introduce a fresh [CLS] token in the second round. We additionally evaluated the use of KV-cache reuse [vaswani2017attention], originally introduced for generative models, and include the results in Table 4. Ultimately, we find that the simplest strategy, Final-Layer Token Recycling, which reuses the features from the last layer of the first round, on average, achieves the best performance. This indicates that the output of the first round already captures sufficient high-level information to guide the second round effectively, while being comparatively simple to implement. It is important to note that except the KV-cache experiment, all results were conducted during the early experimental phase of ThinkingViT, and due to training resource constraints, they were not trained with the full joint training strategy described in Section 3.3, but rather with the stochastic training method introduced in [haberer2024hydravit]. As a result, their accuracies are slightly lower than those of the final model reported in Section 4.
@@ Appendix E Ablation of different fusing strategies
Table 5 compares six fusion strategies that recycle the tokens produced in the first round, denoted as $z_{1}$, by incorporating them into the initial patch embeddings of the second round, $\mathcal{E}_{2}$, on ThinkingViT with a configuration of $3H\rightarrow 6H$. Each variant forms the second-round input as a linear blend, where $\alpha$ is a learnable scalar. The strategies differ in how the two embeddings are projected to the same dimensional space. We consider two projection strategies: (I) a parameter-free approach that either repeats the embedding or pads the lower-dimensional embeddings with zeros, and (II) a learnable linear projection. Empirically, the learnable projection achieves the best overall accuracy by having the highest accuracy in the second round while maintaining performance comparable to the strongest model configuration in the first round. In addition, we evaluate initializing $\alpha=1$ and observe that initializing with zero consistently leads to the best performance. This initialization allows the model to begin training without relying on information from the first round, thereby enabling it to learn the optimal degree of reuse. For example, in ThinkingViT configured with 3H $\rightarrow$ 6H heads, the learned value of $\alpha$ converges to 0.19-0.19.
@@ Appendix G Comparing and combining ThinkingViT with other adaptive baselines
We note that ThinkingViT performs routing at the image level, which is orthogonal to token-level pruning [han2023dynamic, wang2021not, wang2020glance], and thus can be combined with such approaches to provide additional flexibility. To demonstrate this, we incorporate the token pruning mechanism of DynamicViT [rao2021dynamicvit] into the second round of ThinkingViT 3H $\rightarrow$ 6H, applying a pruning ratio of 0.8. As shown in Fig. 14, integrating DynamicViT further improves both efficiency and accuracy, indicating that ThinkingViT 3H $\rightarrow$ 6H is compatible with and complementary to existing token pruning approaches.
@@
Appendix I ImageNet-21K results
To evaluate ThinkingViT on a larger-scale classification task, we train ThinkingViT, MatFormer, HydraViT, DeiT-Tiny, and DeiT-Small on ImageNet-21K [ILSVRC15] as follows:
We use the shuffled dataset imagenet-w21-wds  [^1] https://huggingface.co/datasets/timm/imagenet-w21-wds with the first 12,741,248 images as the training split and the remaining 411,008 images as the validation split. All models are trained from scratch for 150 epochs with a global batch size of 2048.
Figure 16 reports accuracy versus GMACs on the validation split. In the low-compute regime, i.e., below 3.5 GMACs, all nested methods perform comparably. However, as the compute budget increases, ThinkingViT pulls ahead: at roughly 4.6 GMACs, ThinkingViT reaches 41.9% accuracy, outperforming both MatFormer at 39.8% and HydraViT at 39.4% by more than 2 percentage points. Moreover, ThinkingViT offers a higher accuracy ceiling of 42.4% by leveraging its second inference round for difficult samples, matching standalone DeiT-Small at 42.3% while providing a continuous range of cheaper operating points through entropy-based early exit.
These results show that the benefits of ThinkingViT generalize beyond ImageNet-1K to large-scale settings with an order of magnitude more classes.
Figure 16: Accuracy versus GMACs on ImageNet-21K. All models are trained from scratch for 150 epochs. While nested methods perform similarly below 3.5 GMACS, ThinkingViT achieves a higher accuracy ceiling. Stars denote standalone DeiT-Tiny (T) and DeiT-Small (S).
Appendix J Limitations
Training Overhead: Compared to training multiple standalone models, nested models like ThinkingViT incur higher training costs to reach similar performance levels across all stages. This is a known limitation of nested architectures, which share parameters and require joint optimization to maintain accuracy at varying compute levels. In our case, ThinkingViT 3→6 trains DeiT-T and DeiT-S with a two-stage forward and single backward pass, taking  20 min/epoch, comparable to training DeiT-T (7 min) and DeiT-S (13 min) separately. Therefore training ThinkingViT takes as long as training DeiT-T and DeiT-S for 300 epochs each.
Jumping Too Far Reduces Effectiveness:
When the model transitions from a very small subset of attention heads (e.g., 3H) directly to a full configuration (e.g., 12H) in the second stage, performance suffers compared to using an intermediate stage (e.g., 9H) instead. This suggests that overly aggressive compute expansion can disrupt representational continuity, emphasizing the importance of smooth progression in staged inference.
Appendix K Batch inference
@@ Appendix I ImageNet-21K results
Appendix L Effect of attention-head expansion and loss weighting on multi-round thinking